OpenStudy (anonymous):

729 to the power -1/6

OpenStudy (anonymous):

1/3

OpenStudy (aum):

Write 729 as 3^6 and simplify.

OpenStudy (aum):

\[ 729^{-1/6} = (3^6)^{-1/6} = 3^{-1} = 1/3 \]